Let {an} sequence is defined as a1=3, an+1=1/5(an)


converges to zero. Prove that




1
Expert's answer
2022-01-11T00:03:47-0500

Solution:

"a_1=3, a_{n+1}=\dfrac15(a_n) 
\\ \Rightarrow a_{2}=\dfrac15(a_1)= \dfrac15(3)=\dfrac35
\\ \Rightarrow a_{3}=\dfrac15(a_2)= \dfrac15(\dfrac35)=\dfrac3{25}
\\ \Rightarrow a_{4}=\dfrac15(a_3)= \dfrac15(\dfrac3{25})=\dfrac3{125}"

Thus, "a_{n+1}=\dfrac{3}{5^n}"

Now, "\lim_{n\rightarrow \infty}a_{n+1}=\lim_{n\rightarrow \infty}\dfrac{3}{5^n}"

"=\dfrac{3}{\infty}
\\=0"